resistor in the following circuit using the Node Solve for the current lo through the 2 Voltage Method. 41 2 M 12 V
Answer #1

1. We need Is through an. 4n Vio - Vi - ar vi er 1 V2 I. = 2 2 Io So to find the corrent (10) 12v 6 2.0 through an Resistor.substitute Vy = 8 in Equation o 7(8) - V2 = 36 :: V2 = 56-36 = 20V. the need Iosu - 8 = 4 Amp D. Io = 4A. E A .